### v1

# Mitigating Nonlinear Systematics in Weak Lensing Surveys II: Stability and Diagnostics with Intrinsic Alignment

Abstract
The Bernardeau–Nishimichi–Taruya (BNT) transform provides a powerful framework for analysing tomographic cosmic shear data by improving the localization of shear correlations in physical scale. It operates by performing a linear combination of the shear data vector in $\ell$-space, yielding a transformed vector that is better localized in both redshift and $k$-space. BNT is particularly useful for estimating cosmological parameters while minimizing the impact of poorly understood nonlinear physics, without discarding large amounts of information as is typically done with simple scale cuts. In our previous work, we showed that BNT outperforms traditional weak-lensing analyses; however, that study did not include intrinsic alignments (IA). In the present work, we assess the robustness of our BNT-based $k$-cut framework in the presence of realistic IA models. We consider two cases: (i) when the assumed IA model used in sampling is close to, but not identical to, the true one, and (ii) when the assumed IA model is significantly mis-specified. In the first case, the $k$-cut framework yields precise and unbiased $S_{8}$ constraints even with limited knowledge of large-scale modes. In the second, the BNT transform serves as a powerful diagnostic tool, revealing internal inconsistencies in the assumed IA model across $k$- and redshift-space, indicating that the corresponding cosmological constraints cannot be trusted until a more accurate IA model is adopted.
preprint: APS/123-QED
I Introduction
Weak gravitational lensing as a probe of dark matter is a core science objective of Stage IV imaging surveys (e.g., Euclid [1], Rubin [2], Xuntian [3]). The substantial gain in statistical power provides a powerful means to identify and reduce residual systematics, whether instrumental (e.g., Point Spread Function characterization and correction) or theoretical (e.g., modeling of the nonlinear matter power spectrum). At the precision expected from Stage IV surveys, even subtle biases arising from incomplete models, instrumental artifacts, or intrinsic astrophysical correlations could dominate the error budget and bias cosmological inference. Mitigating these biases is now a defining challenge for weak-lensing cosmology.
Weak gravitational lensing is fundamentally a two-dimensional probe: for a given source redshift, it measures anisotropies in the mass distribution projected along the line of sight from the sources to the observer. This makes it challenging to identify biases in the nonlinear three-dimensional modeling of the matter power spectrum, because only two-dimensional projections are observed. Consequently, the usual approach is to apply scale cuts to the WL data vector, at the cost of significant information loss. However, [4] showed that a certain linear combination of weak-lensing observables can recover some of the three-dimensional information. This is possible because, for a lens at a given redshift, the redshift scaling of the WL effect is purely geometric and independent of the matter distribution—a property that underlies nulling techniques [5, 6]. The specific linear combination that achieves this, developed in [4], is called the Bernardeau–Nishimichi–Taruya (BNT) transform. In our previous work [7] (Paper I), we adapted and tested this transform as a targeted solution for mitigating nonlinear systematics.
We showed that a BNT-transformed data vector, combined with a well-defined angular scale filtering ($\ell$-cut in harmonic space), can effectively operate as a $k$-cut, enabling more robust handling of uncertainties in the nonlinear modeling of the three-dimensional matter power spectrum. However, in Paper I we did not account for intrinsic alignment (IA), a known contaminant of the WL signal that can mimic lensing by correlating galaxy shapes over large distances. In this paper, we revisit a number of results obtained in paper I and reevaluate the performance of the BNT transform when IA is taken into account.
The paper is organised as follows. In Section II, we review the modelling of cosmic shear and intrinsic alignment signals, along with their corresponding BNT transforms. In Section III, we describe our methodology for calculating physical and redshift scale cuts using BNT, and outline our posterior sampling procedure. In Section IV, we present forecasts of parameter constraints using both BNT and non-BNT approaches, applied to Euclid-like mock data with non-zero intrinsic alignment, and demonstrate the tests and consistency checks introduced above. Finally, in Section V, we summarise our findings and offer concluding remarks.
II Theory
II.1 Cosmic Shear
We consider a cosmological survey with measured photometric redshifts for galaxies, enabling tomographic binning with source-galaxy distributions $n_{i}(\chi)$ for the $i$th bin. The two-dimensional cosmic-shear power spectrum $C^{i,j}(\ell)$ between tomographic bins $(i,j)$ is given by the line-of-sight projection of the non-linear three-dimensional matter power spectrum $P_{\rm nl}(k;z)$, as follows [8]:
$C^{i,j}(\ell)=\int\dfrac{d\chi}{\chi^{2}}W^{i}_{\gamma}(\chi)W^{j}_{\gamma}(\chi)P_{\rm nl}\left(k=\dfrac{\ell+1/2}{\chi};z(\chi)\right),$ (1)
where $\chi$ is the radial comoving distance at redshift $z$, and $W_{\gamma}^{i}(\chi)$ is the lensing kernel computed by the source redshift distribution of the $i^{th}$ tomographic bin:
$W_{\gamma}^{i}(\chi)=\dfrac{\Omega_{m}^{2}H_{0}^{4}}{c^{2}}\int\text{d}\chi^{\prime}\dfrac{n_{i}(\chi^{\prime})}{a(\chi)}\dfrac{(\chi^{\prime}-\chi)\chi}{\chi^{\prime}},$ (2)
where $a(\chi)$ is the scale factor. We have assumed a flat Universe $K=0$, for non-flat geometry, the radial comoving distance $\chi$ has to be replaced by the comoving angular diameter distance $f_{K}(\chi)$ in Eq. 2:
$f_{K}(\chi)=\begin{cases}\dfrac{1}{\sqrt{K}}\;\sin\!\big(\sqrt{K}\,\chi\big),&K>0\ \text{(closed)},\\[6.0pt] \chi,&K=0\ \text{(flat)},\\[6.0pt] \dfrac{1}{\sqrt{-K}}\;\sinh\!\big(\sqrt{-K}\,\chi\big),&K<0\ \text{(open)}.\end{cases}$
II.2 Intrinsic Alignment
In addition to the coherent distortions induced by gravitational lensing described by Eq. 1, intrinsic alignments (IA) generate additional angular correlations among galaxy shapes—both within a given tomographic bin, through local tidal interactions, and across widely separated bins, via long-range correlations. These intrinsic correlations can mimic or bias the lensing signal, making IA a significant astrophysical systematic in weak-lensing analyses [9].
One of the simplest IA models is the Non-linear Alignment (NLA) model, which extends the linear alignment framework [10, 9] by replacing the linear matter power spectrum with its nonlinear counterpart [11]. It assumes that intrinsic alignments arise from a tidal interaction that linearly couples galaxy shapes to the local large-scale gravitational field, while nonlinear clustering effects are incorporated through the use of $P_{\mathrm{nl}}(k)$. The NLA model can be described by a redshift-dependent linear bias $C_{1}(z)$ applied to the nonlinear matter power spectrum:
$P_{\mathrm{GG}}(k,z)$ $=$ $P_{\mathrm{nl}}(k;z)$
$P_{\mathrm{GI}}(k,z)$ $=$ $C_{1}(z)\,P_{\mathrm{nl}}(k;z)$
$P_{\mathrm{II}}(k,z)$ $=$ $C_{1}^{2}(z)\,P_{\mathrm{nl}}(k,z)$ (3)
where the tidal alignment bias $C_{1}(z)$ is always parameterised as a redshift-dependent function involving a global IA scaling amplitude $A_{\mathrm{TA}}$ and a power-law evolution index $\eta_{\mathrm{TA}}$:
$C_{1}(z)\equiv-A_{\mathrm{TA}}\,\tilde{C}_{1}\,\frac{\rho_{\mathrm{crit}}\,\Omega_{m}}{D(z)}\left(\frac{1+z}{1+z_{0}}\right)^{\eta_{\mathrm{TA}}}$ (4)
where $\rho_{\mathrm{crit}}$ is the critical density of the universe, $D(z)$ is the structure growth function, $z_{0}$ is a pivot redshift often chose to be $z_{0}=0.62$ [12], and $\tilde{C}_{1}=5\times 10^{-14}\,h^{-2}\mathrm{M}_{\odot}^{-1}\mathrm{Mpc}^{3}$ is a normalisation constant.
There is more than one approach for addressing the redshift dependence of the intrinsic alignment. A commonly used alternative to the $\eta_{\mathrm{TA}}$ parametrization involves the $B_{\mathrm{TA}}$ parameter in a so-called NLA-z model [13]:
$C_{1}(z)=-\tilde{C}_{1}\,\frac{\rho_{\mathrm{crit}}\,\Omega_{m}}{D(z)}\left[A_{\mathrm{TA}}+B_{\mathrm{TA}}\left(\frac{\langle a(z)\rangle_{i}}{a_{\mathrm{piv}}}-1\right)\right],$ (5)
where $\langle a(z)\rangle_{i}$ is the $n(z)$-weighted average scale factor in tomographic bin $i$, and $a_{\mathrm{piv}}=0.769$ is the pivotal scale factor. Note that, unlike the $\eta_{\mathrm{TA}}$ parameterisation of NLA model, this NLA-z model can change sign as a function of redshift.
Another frequently-used model is the Tidal Alignment and Tidal Torquing (TATT) model [14], which extends beyond the NLA framework by incorporating both linear (Tidal Alignment, TA) and quadratic (Tidal Torquing, TT) responses of galaxy shapes to the tidal field. Developed to capture intrinsic alignments of both early- and late-type galaxies within a unified framework, the TATT model introduces additional degrees of freedom to describe more complex IA behaviour. However, even if TATT offers greater physical flexibility, its additional parameters are not yet tightly constrained by current data, leading to substantial degeneracies and potential biases when applied to lensing surveys [15].
In the TATT model, the intrinsic shape of a galaxy can be written as an expansion of the tidal field tensor $s_{ij}$ where [14]:
$\gamma_{ij}^{\mathrm{IA}}=C_{\mathrm{1}}s_{ij}+C_{1\delta}(\delta\times s_{ij})+C_{2}\left[\sum_{k=0}^{2}s_{ik}s_{kj}-\frac{1}{3}\delta_{ij}s^{2}\right]+$... (6)
where $\delta$ is the matter overdensity, $\delta_{ij}$ is the kronecker delta, and tidal field tensor $s_{ij}$ is:
$s_{ij}({\bm{k}})=\left(\frac{k_{i}k_{j}}{k^{2}}-\frac{1}{3}\delta_{ij}\right)\delta({\bm{k}})$ (7)
with $k^{2}=k_{1}^{2}+k_{2}^{2}+k_{3}^{2}$, and $k$ is the modulus of the 3-dimensional wavevector $\bm{k}$. For $C_{1\delta}$ and $C_{2}$, a similar description like Eq. 4 can be used with some minor modifications:
$C_{1\delta}(z)$ $=$ $b_{\mathrm{TA}}C_{1}(z)$
$C_{2}(z)$ $=$ $-5A_{\mathrm{TT}}\,\tilde{C}_{1}\,\frac{\rho_{\mathrm{crit}}\,\Omega_{m}}{D^{2}(z)}\left(\frac{1+z}{1+z_{0}}\right)^{\eta_{\mathrm{TT}}}$ (8)
where $A_{\mathrm{TT}}$ is the Tidal Torquing scaling amplitude and $b_{\mathrm{TA}}>0$ is the density weighting bias. To avoid introducing too many equations, please refer to the Section III. of [14] for the perturbation theory-based prescription to compute $P_{\mathrm{GI}}(k)$ and $P_{\mathrm{II}}(k)$ under the TATT model. It is worth mentioning that, for the TATT model, Eq. 3 cannot be applied. This is a consequence of Eq. 6 which implies that the redshift and $k$ dependence can no longer be factorized.
After obtaining the power spectra $P_{\mathrm{GG}}$, $P_{\mathrm{GI}}$ and $P_{\mathrm{II}}$, we can project it to the 2-D celestial sphere with a similar procedure mentioned in the Eq. 1:
$C_{\gamma^{i},\gamma^{j}}(\ell)$ $=$ $\int\dfrac{d\chi}{\chi^{2}}W^{i}_{\gamma}(\chi)W^{j}_{\gamma}(\chi)P_{\mathrm{GG}}(k;z)$ (9)
$C_{\gamma^{i},I^{j}}(\ell)$ $=$ $\int\dfrac{d\chi}{\chi^{2}}W^{i}_{\gamma}(\chi)W^{j}_{N}(\chi)P_{\mathrm{GI}}(k;z)$ (10)
$C_{I^{i},\gamma^{j}}(\ell)$ $=$ $\int\dfrac{d\chi}{\chi^{2}}W^{i}_{N}(\chi)W^{j}_{\gamma}(\chi)P_{\mathrm{GI}}(k;z)$ (11)
$C_{I^{i},I^{j}}(\ell)$ $=$ $\int\dfrac{d\chi}{\chi^{2}}W^{i}_{N}(\chi)W^{j}_{N}(\chi)P_{\mathrm{II}}(k;z)$ (12)
and
$C^{i,j}\equiv C_{\gamma^{i},\gamma^{j}}+C_{\gamma^{i},I^{j}}+C_{I^{i},\gamma^{j}}+C_{I^{i},I^{j}},$ (13)
where $W^{i}_{N}(\chi)$ is the number density projection kernel:
$W^{i}_{N}(\chi)=b^{i}\frac{H(\chi)}{c}\,n_{i}(z(\chi)),$ (14)
and $b^{i}$ being the linear galaxy bias of ith tomographic bin. In this paper, in order to keep the focus on the intrinsic alignment, we set $b^{i}=1$. For an effective description of the intrinsic alignment of its variation compared to the standard NLA model (Eq. 4), we also define an effective tidal alignment amplitude $A_{\mathrm{TA,eff}}$:
$A_{\mathrm{TA,eff}}(k,z)$ $\equiv$ $\dfrac{D(z)}{\tilde{C}_{1}\rho_{\mathrm{crit}}\Omega_{m}}\sqrt{\dfrac{P_{\mathrm{II}}(k,z)}{P_{\mathrm{GG}}(k,z)}}$
$=$ $-\dfrac{D(z)}{\tilde{C}_{1}\rho_{\mathrm{crit}}\Omega_{m}}C_{1}(z)$
II.3 BNT Transform
The BNT transform is a linear combination of the data vector $C^{i,j}(\ell)$ into a new one $\hat{C}^{a,b}(\ell)$, where $a$ and $b$ represent a linear combination of the original tomographic redshift bins $i$ and $j$. Throughout the paper, we will keep this ‘hat’ notation $\hat{}$ for all BNT transformed quantities, and reserve the use of $(i,j)$ for the original, noBNT tomographic bins and $(a,b)$ for the BNT tomographic bins.
The BNT transform matrix $p^{a}_{i}$ is defined as the following. Two normalisation numbers, $n_{i}^{0}$ and $n_{i}^{1}$, are first calculated for each tomographic bin $i$:
$n_{i}^{0}$ $=$ $\int\text{d}\chi\;n_{i}(\chi)$ (16)
$n_{i}^{1}$ $=$ $\int\text{d}\chi\;\dfrac{n_{i}(\chi)}{\chi}$ (17)
The transformation matrix elements $p^{a}_{i}$ are then constructed from the following algebraic equations:
$\sum_{i=a-2}^{a}$ $p^{a}_{i}n_{i}^{0}=0$ (18)
$\sum_{i=a-2}^{a}$ $p^{a}_{i}n_{i}^{1}=0,$
with $p^{a}_{i}=0$ when $i\notin\{a-2,a-1,a\}$, $p_{1}^{1}=p_{2}^{2}=1$, and $p_{2}^{1}=-1$. The matrix $p^{a}_{i}$ is hence a $n_{\rm T}\times n_{\rm T}$ square matrix, where $n_{\rm T}$ is the number of tomographic bins, with non-zero elements only in a diagonal band of a width of three. Fig. 1 shows the BNT transformation matrix for our fiducial cosmology: the columns represent the original tomographic bin entries and the rows represent the BNT tomographic bin entries. One can see that for all rows, except the first, the sum of the matrix elements is always zero. This is a feature of the transform construction, which is built from the algebraic nulling constraints in Eq. 18.
Figure 1: The BNT transform matrix $p^{a}_{i}$ calculated with our fiducial cosmology and redshift distribution.
Unlike Paper I [7], where intrinsic alignment was not included in the BNT modelling, we now apply the BNT matrix to the full data vector defined in Eq. 13, including the intrinsic-alignment terms. This is the only consistent approach, as the observer cannot separate the lensing component $C_{\gamma^{i},\gamma^{j}}$ from the other contributions:
$\hat{C}^{a,b}(\ell)=p^{a}_{i}p^{b}_{j}C^{i,j}(\ell)$ (19)
It is worth mentioning that, for the NLA model, combining Eq. 3 and Eqs. 9-12 leads to a compact form for the BNT data vector $\hat{C}^{a,b}(\ell)$:
$\hat{C}^{a,b}(\ell)=\int\dfrac{d\chi}{\chi^{2}}\hat{W}^{a}_{\rm eff}(\chi)\hat{W}^{b}_{\rm eff}(\chi)P_{\mathrm{GG}}(k;z),$ (20)
where we have defined:
$\hat{W}^{a}_{\mathrm{eff}}(\chi)\hat{W}^{b}_{\mathrm{eff}}(\chi)$ $\equiv$ $\hat{W}_{\gamma}^{a}(\chi)\hat{W}_{\gamma}^{b}(\chi)+\hat{W}_{\gamma}^{a}(\chi)\hat{W}_{I}^{b}(\chi)$ (21)
$+$ $\hat{W}_{I}^{a}(\chi)\hat{W}_{\gamma}^{b}(\chi)+\hat{W}^{a}_{I}(\chi)\hat{W}^{b}_{I}(\chi),$
with:
$\hat{W}_{\gamma}^{a}(\chi)$ $\equiv$ $\sum_{i=1}^{n_{\rm T}}p^{a}_{i}W_{\gamma}^{i}(\chi)$
$\hat{W}^{a}_{I}(\chi)$ $\equiv$ $\sum_{i=1}^{n_{T}}p^{a}_{i}\left[C^{i}_{1}(\chi)W_{N}^{i}(\chi)\right].$ (22)
With this definition, $\hat{W}_{\gamma}^{a}(\chi)$ is the BNT-transformed lensing kernel and $\hat{W}^{a}_{I}(\chi)$ is the BNT-transformed intrinsic-alignment kernel. It is important to note that, by construction, the BNT lensing kernel is localized in redshift, whereas this is not necessarily true for the BNT intrinsic-alignment kernel. The remainder of this work quantifies how intrinsic alignments affect the ability of BNT to separate angular from physical scales in the lensing term. For TATT in particular, the factorization in Eq. 20 is not applicable due to the $k$ dependence in the IA kernel.
Figure 2: Panel 1 of 4: BNT Lensing projection weight $\hat{W}_{\gamma}^{2}(\chi)/\chi^{2}$ for the shear-shear term. The original BNT projection kernel. Panel 2 of 4: BNT Lensing projection weight $\hat{W}_{\gamma}(\chi)\hat{W}_{I}(\chi)/\chi^{2}$ for the shear-shape term with $W_{I}(\chi)=C_{1}(\chi)W_{N}(\chi)$ under the NLA model with $A_{\mathrm{TA}}=1.71$. Panel 3 of 4: BNT Lensing projection weight $\hat{W}_{I}^{2}(\chi)/\chi^{2}$ for the shape-shape term with $W_{I}(\chi)=C_{1}(\chi)W_{N}(\chi)$ under the NLA model with $A_{\mathrm{TA}}=1.71$. Panel 4 of 4: The total projection weight $W^{2}_{\mathrm{eff}}(\chi)=\hat{W}_{\gamma}^{2}(\chi)+2\hat{W}_{\gamma}(\chi)\hat{W}_{I}(\chi)+\hat{W}_{I}^{2}(\chi)$ for the shear and shape included total $C_{\ell}$.
In the rest of this work, we will use Euclid-like mock with ten redshift bins [16]. Figure 2 displays the three possible kernel combinations as functions of $\chi$, alongside the effective kernel $\chi^{-2}\hat{W}^{2}_{\mathrm{eff}}(\chi)$. The BNT intrinsic-alignment kernel is less localized in redshift than the BNT lensing kernel, yet it remains more localized than the original noBNT lensing kernel.
III Methodology
III.1 Scale Cuts
Figure 3: Ratio of the data vector with $k_{\rm cut}$ to the full data vector (defined in Eq. 23) for the noBNT case $\mathfrak{R}^{(i,j)}(\ell;k_{\rm cuts},\Pi^{0})$ (upper-right triangle) and the BNT case $\hat{\mathfrak{R}}^{(i,j)}(\ell;k_{\rm cuts},\Pi^{0})$ (lower-left triangle). Each panel shows the ratio for the combination of tomographic bins indicated by $(i,j)$ (for the noBNT case) and $(a,b)$ (for the BNT case). For each tomographic bin, the curves are used to define which cut in $\ell$ corresponds to the constraint that modes $k>k_{\rm high}$ and $k<k_{\rm low}$ should not bias the data vector at a level exceeding $\mathcal{T}_{\mathrm{FD}}$ for $\ell_{\rm low}^{{(i,j)}}<\ell<\ell_{\rm high}^{{(i,j)}}$ for noBNT (resp. $\hat{\ell}_{\rm low}^{{(a,b)}}<\ell<\hat{\ell}_{\rm high}^{{(a,b)}}$ for BNT). All plots are given for $k_{\mathrm{cuts}}=(0.1\;{\rm Mpc^{-1}},1.0\;{\rm Mpc^{-1}})$, and the horizontal bright band indicates the fractional threshold $\mathcal{T}_{\mathrm{FD}}=\pm 0.1$. For each tomographic bin combination, the cut in $\ell$ is shown as the boundary of the cyan and green regions.
We adopt the same approach as in paper I, where we determined which $\ell$-scale cut $\ell<\ell_{\rm cut}$ corresponds to a physical $k$-cut $k<k_{\rm cut}$ for a given accuracy. The difference here is that we now consider a domain cut, $k_{\mathrm{cuts}}=[k_{\mathrm{low}},k_{\mathrm{high}}]$, and attempt to match $k_{\mathrm{cuts}}$ to a corresponding range in $\ell$-space such that the resulting data vector remains accurate to a specified precision within this $\ell$-space interval. The relationship between $k$ and $\ell$ cuts for both of the BNT and noBNT data vectors will be estimated using these dimensionless ratios:
$\hat{\mathfrak{R}}^{(a,b)}(\ell;k_{\rm cuts},\Pi^{0})$ $\equiv$ $\frac{\hat{\mathcal{C}}^{a,b}_{[k_{\mathrm{low}},k_{\mathrm{high}}]}(\ell;\Pi^{0})}{\hat{\mathcal{C}}^{a,b}_{[k_{\mathrm{min}},k_{\mathrm{max}}]}(\ell;\Pi^{0})}$
${\mathfrak{R}}^{(i,j)}(\ell;k_{\rm cuts},\Pi^{0})$ $\equiv$ $\frac{{\cal C}^{i,j}_{[k_{\mathrm{low}},k_{\mathrm{high}}]}(\ell;\Pi^{0})}{{\cal C}^{i,j}_{[k_{\mathrm{min}},k_{\mathrm{max}}]}(\ell;\Pi^{0})}$ (23)
where $k_{\mathrm{min}}=0.0025$ Mpc-1 and $k_{\mathrm{max}}=100$ Mpc-1 is the entire computational domain in $k$-space. As noted in Paper I, the BNT transform does not need to be computed with the true cosmology. Actually, searching for the true cosmology for which nulling would be exact can itself be exploited ([17, 18]).
Following Paper I, we also define a fractional difference threshold $\mathcal{T}_{\mathrm{FD}}$ to characterise how a $\ell$-cut is defined based on the tolerance region with $\hat{\mathfrak{R}}=1\pm\mathcal{T}_{\mathrm{FD}}$. In order to include the impact of both $k_{\mathrm{cuts}}$ and $\mathcal{T}_{\mathrm{FD}}$, we use the same continuous Boolean weight function $\mathcal{V}(k_{\mathrm{cuts}},\mathcal{T}_{\mathrm{FD}})$ as the paper I:
${\mathcal{V}}(k_{\mathrm{cuts}},\mathcal{T}_{\mathrm{FD}})=\begin{cases}1,|\mathfrak{R}(k_{\rm cuts},\Pi^{0})-1|\leq\mathcal{T}_{\mathrm{FD}}\\ 0,|\mathfrak{R}(k_{\rm cuts},\Pi^{0})-1|>\mathcal{T}_{\mathrm{FD}},\end{cases}$ (24)
where $\mathfrak{R}(k_{\rm cuts},\Pi^{0})$ is the dimensionless power spectra ratio computed by the entire tomographic data vector, and can be applicable for the cases of noBNT and BNT. This Boolean weight function is used in the likelihood computation:
$\Delta{\cal D}^{*}$ $\equiv$ $\mathcal{V}(k_{\mathrm{cuts}},\mathcal{T}_{\mathrm{FD}})\left({\cal C}(\Pi)-{\cal C}(\Pi^{0})\right)$
$\mathbb{C^{*}}$ $\equiv$ $\mathcal{V}(k_{\mathrm{cuts}},\mathcal{T}_{\mathrm{FD}})\cdot\mathbb{C}\cdot\mathcal{V}(k_{\mathrm{cuts}},\mathcal{T}_{\mathrm{FD}})^{{}^{T}}$
$\chi^{2}(\Pi;k_{\mathrm{cuts}},\mathcal{T}_{\mathrm{FD}})$ $=$ $\Delta{{\cal D}^{*}}^{T}\cdot(\mathbb{C^{*}})^{-1}\cdot\Delta{\cal D}^{*}$ (25)
where ${\cal C}(\Pi)$ is the entire tomographic data vector that is organized like:
${\cal C}_{[k_{1},k_{2}]}(\Pi)=\begin{pmatrix}\begin{array}[]{cc}\begin{bmatrix}C^{1,1}_{[k_{1},k_{2}]}(\ell;\Pi)\\ \end{bmatrix}{\rm all}~\ell^{\prime}s\\ \begin{array}[]{c}\end{array}\\ \begin{bmatrix}C^{1,2}_{[k_{1},k_{2}]}(\ell;\Pi)\\ \end{bmatrix}{\rm all}~\ell^{\prime}s\\ \begin{array}[]{c}\end{array}\\ \begin{bmatrix}C^{1,3}_{[k_{1},k_{2}]}(\ell;\Pi)\\ \end{bmatrix}{\rm all}~\ell^{\prime}s\\ ...&\\ \begin{array}[]{c}\end{array}\\ \begin{bmatrix}C^{9,10}_{[k_{1},k_{2}]}(\ell;\Pi)\\ \end{bmatrix}{\rm all}~\ell^{\prime}s\\ \begin{array}[]{c}\end{array}\\ \begin{bmatrix}C^{10,10}_{[k_{1},k_{2}]}(\ell;\Pi)\\ \end{bmatrix}{\rm all}~\ell^{\prime}s\\ \end{array}\end{pmatrix}$. (26)
The $C^{i,j}_{\ell}$ or $\hat{C}^{a,b}_{\ell}$ difference is used for distinguish $i,j$ (noBNT) or $a,b$ (BNT) tomographic bins, and $\mathbb{C}$ is the corresponding covariance matrix.
Figure 3 shows the impact of a particular $k$-cut on the angular power spectrum $C_{\ell}$ for different tomographic combinations. It results in a biased $C_{\ell}$ for a range of $\ell$ which depends on the tomographic bin. As expected, the noBNT $C_{\ell}$ is significantly more biased than the BNT one. In order to quantify this bias, we define, for BNT, $\hat{\ell}^{(a,b)}_{\rm cuts}=(\hat{\ell}^{(a,b)}_{\rm low},\hat{\ell}^{{(a,b)}}_{\rm high})$ (and $\ell^{(i,j)}_{\rm cuts}$ for noBNT) as the range of $\ell$ where the bias does not exceed $\mathcal{T}_{\mathrm{FD}}$, i.e. where $\hat{\mathfrak{R}}^{(a,b)}(\ell\in\hat{\ell}^{a,b}_{\rm cuts};k_{\rm cuts},\Pi^{0})\in(1-\mathcal{T}_{\mathrm{FD}},1+\mathcal{T}_{\mathrm{FD}})$.
[…]
III.2 Tomographic redshift bins cuts
In addition to applying $k$-band cuts, we can also perform the measurement within a subset of redshift bins. The motivation for doing these measurements in selected redshift bins is to verify the consistency of the cosmological parameter inference across bins. This is particularly relevant here because the redshift dependences of IA and lensing differ. Using the Euclid-like mock with ten redshift bins, we partition the lensing kernels into three groups: $\mathbb{Z}_{1,2,3}$ for the first three kernels $\{1,2,3\}$, $\mathbb{Z}_{4,5,6}$ for kernels $\{4,5,6\}$, and $\mathbb{Z}_{7,8,9,10}$ for kernels $\{7,8,9,10\}$. This partition is applied to both the non-BNT and BNT data vectors. For BNT, it is applied to the BNT-transformed lensing kernels $(a,b)$ in $C^{a,b}(\ell)$, whereas for the non-BNT case it is applied to the non-BNT lensing kernels $(i,j)$ in $C^{i,j}(\ell)$.
This binning strategy is designed to assess any possible inconsistency between redshift bins. For each lensing-kernel group, we include in the data vector all auto- and cross-correlations for which both indices e.g. $(a,b)$ for BNT, belong to the same group, and exclude any cross-correlations whose indices come from different groups. For example, under this scheme, the spectrum $C^{2,3}(\ell)$ is included in $\mathbb{Z}_{1,2,3}$, while $C^{3,4}(\ell)$ is excluded everywhere because its indices belong to different groups ($\mathbb{Z}_{1,2,3}$ and $\mathbb{Z}_{4,5,6}$).
III.3 Analysis setup
Table 1: The prior of all the parameters, except for IA parameters, used in our Nautilus sampling. $\mathcal{U}(a,b)$ is a flat prior between $a$ and $b$, and $\mathcal{N}(c,d)$ is a Gaussian prior that centred at $c$ with standard deviation $d$.
Parameters Fiducial Value Priors
$\Omega_{m}$ $0.2905$ $\mathcal{U}(0.1,0.6)$
$10^{9}A_{s}$ $2.1868$ $\mathcal{U}(1.5,5.0)$
$\Omega_{b}$ $0.0473$ $\mathcal{U}(0.03,0.07)$
$n_{s}$ $0.9690$ $\mathcal{U}(0.92,1.02)$
$h$ $0.6898$ $\mathcal{U}(0.55,0.85)$
$A_{\mathrm{B}}$ $3.1300$ $\mathcal{U}(1.0,6.0)$
$\sigma_{8}$ $0.8256$ Derived Parameter
$S_{8}$ $0.8124$ Derived Parameter
The analysis presented here extends our Paper I study by incorporating intrinsic alignment and applying a narrower $k$-cut, while retaining the same overall objective: to compare the BNT and non-BNT cases in terms of (1) the precision of the cosmological parameter measurements, and (2) the bias of the posterior relative to the input model (including cases where the intrinsic alignment model is invalid). We carry out a series of Bayesian posterior estimations to assess the impact of intrinsic alignment, assuming a ”True Background Cosmology” (TBC) and employing two distinct $P(k)$ models: the canonical HMcode [19] and the non-standard AxionHMcode [20], which accounts for the effects of ultralight axions on the small-scale growth of large-scale structure. In this study, we adopt a fixed axion mass of $m_{\mathrm{axion}}=10^{-24}\,$eV and an axion density of $\Omega_{\mathrm{axion}}=0.15\times\Omega_{m}=0.0436$.
The detailed cosmological parameters for our TBC are listed in Table 1. The fiducial values and priors of these parameters remain identical across all posterior samplings. For the intrinsic alignment component, we consider three different models. The NLA and TATT model parameters follow the fits reported in [21], derived from galaxy correlation functions based on the IllustrisTNG hydrodynamical simulations [22, 23]. In addition, the NLA-z model adopts the best-fit parameters from the KiDS Legacy cosmic shear analysis [24].
Table 2: The prior of all the parameters used in our Nautilus sampling. For table width considerations, we use a simplified notation here $\mathcal{U}_{a,b}=\mathcal{U}(a,b)$ for a flat prior between $a$ and $b$. The single value prior means we fix that parameter to this particular value in that prior. We don’t necessarily sample all parameters below in every chain. If they are not sampled, their values are set to the ’Fiducial Value’ column which indicates the values taken by the parameters for the fiducial cosmology.
Parameters $A_{\mathrm{TA}}$ $\eta_{\mathrm{TA}}$ $B_{\mathrm{TA}}$ $b_{\mathrm{TA}}$ $A_{\mathrm{TT}}$ $\eta_{\mathrm{TT}}$
NLA Model $1.71$ $0.0$ - - - -
TATT Model $1.29$ $0.0$ - $0.21$ $0.32$ $0.0$
NLA-z Model $0.3$ - 3.7-3.7 - - -
NLA Prior $\mathcal{U}_{-6,6}$ $\mathcal{U}_{-5,5}$ - - - -
TATT Prior $\mathcal{U}_{-6,6}$ $\mathcal{U}_{-5,5}$ - $0.0$ $\mathcal{U}_{-6,6}$ $\mathcal{U}_{-5,5}$
NLA-z Prior $\mathcal{U}_{-6,6}$ - $\mathcal{U}_{-6,6}$ - - -
In the first part of our investigation, we assess the stability of the BNT transform under the assumption that we know the IA model precisely. For each of the NLA, TATT, and NLA-z models, we perform posterior sampling using the corresponding prior—i.e., the NLA Prior is run against the fiducial data vector generated with the NLA Model, and likewise for the others. As in Paper I, nonlinear modeling is handled by performing the sampling with HMcode, while the fiducial data vector is generated with AxionHMcode.
In the second part of our analysis, we examine whether the BNT transform can shed light on the validity of assuming that the IA model is accurately known. To this end, we perform sampling under the NLA model while calibrating against fiducial data vectors generated with models that include additional degrees of freedom—specifically, the NLA-z and TATT models. To reduce complexity, both the sampled and fiducial data vectors in this part are generated using the same nonlinear prescription: HMcode. This setup allows us to test the IA model independently of whether the underlying cosmology is assumed to be correct, a possibility enabled by the flexibility of the BNT transform, which permits new and informative manipulations of the data vector.
In addition to the above, the following settings are used for all MCMC runs:
  • We use 50 logarithmically spaced $\ell$-bins, with bin centers ranging from $\ell_{\mathrm{min}}=50$ to $\ell_{\mathrm{max}}=5000$, and adopt the same redshift distribution as in paper I, which also yields the kernels shown in Figure 2.
  • All data vectors used in the posterior sampling are computed with the Core Cosmology Library (CCL) [25], a publicly available and standardized tool for calculating cosmological observables. CCL is developed and maintained by the LSST Dark Energy Science Collaboration (DESC) [26].
  • The covariance matrix is computed using OneCovariance [27], a tool optimized for estimating covariance matrices from two-point statistics in photometric large-scale structure surveys. In our calculation, the statistical noise components are derived assuming an effective survey area of $A_{\mathrm{eff}}=15000\,\mathrm{deg}^{2}$, a galaxy number density per tomographic bin of $n_{\rm gal}=3.0,\mathrm{arcmin}^{-2}$, and a shape noise of $\sigma_{e}=0.3$ [16].
  • Parameter posteriors are obtained with the Nautilus package [28], a highly efficient importance nested-sampling toolkit for Bayesian posterior and evidence estimation.
IV Results
IV.1 BNT and Intrinsic Alignment
We begin by analyzing cases where the intrinsic-alignment (IA) model used to generate the fiducial data vector is also employed in the sampling model, but with different parameter values in each. In the sampling, each parameter is either fixed or varied under the “TATT Prior,” using fiducial data vectors generated with the “TATT Model.” (see Table 2). Note that the $b_{\mathrm{TA}}$ parameter is fixed to different values in these two setups, so the “TATT Prior” does not encompass the “TATT Model.” This is because $b_{\mathrm{TA}}$ is strongly degenerate with $A_{\mathrm{TT}}$ (see Eq. 8). This configuration therefore represents a case where the model is well constrained but not perfectly known—an ideal test of the BNT transform’s stability in the presence of intrinsic-alignment signals. Throughout this section, fiducial data vectors are computed with AxionHMcode, whereas the sampling itself is performed with the standard, no-axion version of HMcode.
Figure 4: Constraints in the $(\Omega_{m},S_{8})$ plane for the noBNT (left panel) and BNT (right panel). The thin solid lines indicate the values of the fiducial cosmology. The power spectrum of all models during the sampling process is calculated using the Mead et al. [19]-version of HMcode, while the fiducial is calculated using the AxionHMcode[20]. The fiducial also included a Samuroff et al. [21]-fit of TATT model from the IllustrisTNG galaxies. Four posterior contours are represented: the Black solid line corresponds to the case where no $k_{\rm cut}$ and no $\mathcal{T}_{\mathrm{FD}}$ are applied. The Lime Green, Dark Red and Cyan contours correspond to $k_{\rm cut}=1.0,0.33,\mathrm{and\ }0.1\;{\rm Mpc^{-1}}$ respectively, and fixed $\mathcal{T}_{\mathrm{FD}}=0.02$. Note that the Cyan contours for noBNT and $k_{\rm cut}=0.1\;{\rm Mpc^{-1}}$ exceeds the plot boundary.
We first examine the joint constraints on $(S_{8},\Omega_{m})$ in a setting where the fiducial nonlinear power spectrum is biased relative to the model used for sampling. As in Paper I, we compare constraints from the BNT and noBNT approaches for various choices of $k_{\rm cut}$, using AxionHMcode with non-zero intrinsic alignments for the fiducial data vector, while the sampling model uses HMcode. In Figure 4, we show the cosmological constraints for different physical scale cuts $k_{\rm cut}$, sampled over a TATT model with $b_{\mathrm{TA}}=0.0$, compared to the fiducial TATT model with $b_{\mathrm{TA}}=0.21$. Notably, this mismatch in $b_{\mathrm{TA}}$ does not introduce any significant additional bias in the cosmological constraints. The bias in $S_{8}$ and $\Omega_{m}$ appears when small-scale information is included, as shown by the shift of contours in Figure 4, which is primarily due to the impact of ultralight axions. However, as we progressively restrict the analysis to larger scales, the $S_{8}$ and $\Omega_{m}$ constraints converge toward the “true background cosmology” indicated by the cross in the contour plot. At $k_{\mathrm{cut}}=0.1\;\mathrm{Mpc}^{-1}$, the recovered cosmology aligns well with the fiducial. These results closely match the result in Paper I, where intrinsic alignment was not included. We conclude that the performance of BNT is not degraded by the presence of intrinsic alignment. By contrast, the noBNT constraints lose most of their constraining power in this regime, a result also similar to Paper I. These results suggest that, even with good but imperfect knowledge of the IA model, the BNT-transform-based $\ell$-cut method remains effective for delivering unbiased cosmological constraints.
Figure 5: Same to the Figure 4, but showing the constraints of $A_{\mathrm{TA}}$ and $A_{\mathrm{TT}}$ instead of $\Omega_{m}$ and $S_{8}$.
Having a different $b_{\mathrm{TA}}$ for the fiducial and sampling models has a significant impact on the IA modeling. This is illustrated in Figure 5 which shows the joint constraints on $A_{\mathrm{TA}}$ and $A_{\mathrm{TT}}$ using the same posterior as in Figure 4. As expected, the noBNT posteriors exhibit a tight and biased constraint with large $k_{\mathrm{cut}}$ values, but give weak constraints when using smaller $k_{\mathrm{cut}}$ values. By contrast, the BNT posterior not only retain its constraining power on cosmological parameters but it also retains its constraining power on $A_{\rm TT}$ and $A_{\rm TA}$ despite a biased $b_{\mathrm{TA}}$ value in the sampling models. Although the location of the maximum posterior in the $(A_{\mathrm{TA}},A_{\mathrm{TT}})$ plane does not move monotonically with changing the scale cut, we recover a well-constrained $A_{\mathrm{TT}}$ value at large-scale $k_{\mathrm{cut}}$ using the BNT transform.
IV.2 Biases induced by Intrinsic Alignment Models
It is possible that the IA model adopted in the sampling is not fully representative of reality. The key question is therefore: to what extent can we trust the cosmological parameter inference, and what diagnostic tools are available to assess the validity of the assumed IA model? An incorrect IA model can, in general, introduce biases either in redshift or in $k$-space. The BNT transform provides access to $k$-space, while redshift-space information can be obtained through $z$-bin grouping, as discussed in Section III.2. In the following, we compare the non-BNT and BNT cases for different $k$-cuts and lensing-kernel groupings.
Figure 6: Constraints in the $(A_{\mathrm{TA}},\eta_{\mathrm{TA}})$ plane for the noBNT (left panel) and BNT (right panel). Note that the fiducial value of $A_{\rm TA}$ is $1.71$ (see Table 2) and therefore outside the x-axis range. The power spectrum of all models for sampling and fiducial are both calculated using the Mead et al. [19]-version of HMcode. The fiducial also included a Wright et al. [24]-fit of NLA-z model from the KiDS-Legacy data, but we fit with an NLA Model against it in our sampling. Four posterior contours are represented: the Black solid line corresponds to the case where no $z_{\rm cuts}$ is applied. The Cyan, Dark Red, and Lime Green contours correspond to the case where we only used the data vector from first three tomographic bins, $4^{\rm th}-6^{\rm th}$ tomographic bins, and last four tomographic bins.
We first investigate the impact of an incorrect IA model in the redshift space by using the NLA-z model from Table 2 for fiducial and NLA model for the sampling. The NLA-z model introduces a different redshift dependence than the standard NLA model, which illustrates the impact of choosing an incorrect IA model. The data vector is split in three tomographic redshift groups: $(i,j)$ or $(a,b)$ for noBNT or BNT, belonging to either $\{1,2,3\}$, $\{4,5,6\}$, or $\{7,8,9,10\}$. For each subset, we exclude cross-correlations between bins belonging to different groups, as explained in Section III.2. Figure 6 shows the resulting constraints on $A_{\mathrm{TA}}$ and $\eta_{\mathrm{TA}}$ derived from the posterior sampling. In both the noBNT and BNT analyses, the resulting contours display substantial inconsistencies across different redshift-bin groupings. Constraints from the first three tomographic bins deviate from those obtained using the full redshift range by more than $5\sigma$ in $\eta_{\mathrm{TA}}$. Likewise, the constraints from the last four tomographic bins differ from those of the first three by over $5\sigma$, while still showing a tension exceeding $4\sigma$ relative to the unsliced case. These findings demonstrate that strong internal inconsistencies can arise when the IA model is incorrect. Redshift grouping therefore provides a useful diagnostic for testing the validity of the assumed IA model.
Figure 7: The effective tidal alignment amplitude $A_{\mathrm{TA,eff}}(z)$ (Eq. II.2) constraints of all three redshift slices derived from the contour of the Figure 6, for both BNT and noBNT cases. The z-values and their errorbars of the data points are derived from the weighted mean and weighted standard deviation, when all galaxies within the $z$-slice (like $\mathbb{Z}_{1,2,3}$) are categorized to be within one single redshift distribution. For the illustration purpose only, the z-value of the noBNT and BNT data points are each shifted by $\pm 0.01$ on the plot, with their $A_{\mathrm{TA,eff}}(z)$ kept to its original constraint value.
Based on the BNT and noBNT constraints shown in Figure 6, we fit an effective tidal alignment amplitude, $A_{\mathrm{TA,eff}}(z)$ (using the definition from Eq.II.2) for each redshift grouping. Figure 7 presents the resulting $A_{\mathrm{TA,eff}}(z)$ compared to the input NLA-z Fiducial model. The centre values and bin width along the $z$-axis are obtained from the $n(z)$-weighted mean and standard deviation, by combining the redshift slices $\mathbb{Z}_{1,2,3}$, $\mathbb{Z}_{4,5,6}$, and $\mathbb{Z}_{7,8,9,10}$ into three effective single redshift distributions. This demonstrates that, for an intrinsic alignment model with explicit redshift dependence, such an approach can serve as a viable method to constrain the detailed redshift evolution of the IA model if the sampling model cannot recover it. For instance, $A_{\mathrm{IA,eff}}(z)$ (Eq. II.2) cannot cross zero with our sampling IA model (Eq. 4), but it can do so with the fiducial IA model (Eq. 5), therefore creating a discrepancy.
Figure 8: Constraints in the $(A_{\mathrm{TA}},\eta_{\mathrm{TA}})$ plane for the noBNT (left panel) and BNT (right panel). The power spectrum of all models for sampling and fiducial are both calculated using the Mead et al. [19]-version of HMcode. The fiducial also included a Samuroff et al. [21]-fit of TATT model from the IllustrisTNG galaxies, but we fit with an NLA Model against it in our sampling. Four posterior contours are represented: the Black solid line corresponds to the case where no $k_{\rm cuts}$ and no $\mathcal{T}_{\mathrm{FD}}$ are applied. The Lime Green, Dark Red and Cyan contours correspond to $k_{\rm cuts}=(0.033,0.33)\;{\rm Mpc^{-1}}$, $(0.1,1.0)\;{\rm Mpc^{-1}}$, and $(0.33,3.3)\;{\rm Mpc^{-1}}$ respectively, and fixed $\mathcal{T}_{\mathrm{FD}}=0.02$. Note that all noBNT contours with $k_{\mathrm{cuts}}$ exceeds the plot boundary. The vertical highlight line on the $A_{\mathrm{TA}}$ gives the NLA fit in the same sample that derives the TATT model we use in the fiducial here.
Similarly, we investigate the impact of $k$-band cuts when the fiducial and sampling IA models differ. In this test, we adopt the $k$-dependent nonlinear TATT model, which introduces additional clustering power at large $k$-scales compared to the conventional NLA model. The fiducial data vector is generated using the TATT model with parameters fitted from the Samuroff et al. [21] study, while sampling is performed with the standard NLA model. Figure 8 shows the constraints on ($\eta_{\rm TA},A_{\rm TA}$) for three $k$-bands: $k_{\mathrm{cuts}}=(0.033,0.33)$, $(0.1,1.0)$, and $(0.33,3.3)\;\mathrm{Mpc}^{-1}$ . Notably, the noBNT analysis shows a substantial loss of constraining power, particularly for the intrinsic-alignment parameter $A_{\mathrm{TA}}$. In contrast, the contours of the BNT analysis are not increasing enormously, showing that the constraining power is preserved across all bands. The constraints on $S_{8}$ remain also consistent across all $k$ bins (not shown). On the other hand, the best fit posterior of $A_{\mathrm{TA}}$ strongly depends on the $k$ bin. There is an increasing systematic shifts from the fiducial value of the TATT model ($A_{\mathrm{TA}}=1.32$) for the low $k$ bin, towards that of the NLA model fit to the same simulated galaxy sample ($A_{\mathrm{TA}}=1.71$). This tension in $A_{\mathrm{TA}}$ between the $(0.033,0.33)$ and $(0.33,3.3)$ bands is much larger than $5\sigma$. This result demonstrates a clear scale-dependent inconsistency: the IA model used in the sampling pipeline fails to capture the scale dependence intentionally introduced in the fiducial – exactly the discrepancy this test is designed to expose. This behaviour is not shared by the noBNT analysis.
Figure 9: The effective tidal alignment amplitude $A_{\mathrm{TA,eff}}(k,z)$ (Eq. II.2) constraints of all three physical scale slices derived from the contour of the Figure 8, for the constraints derived from BNT-transformed data vector at five different redshifts: z=0.0, 0.5, 1.0, 1.5, 2.0z=0.0,\;0.5,\;1.0,\;1.5,\;2.0. The errorbar on $k$-axis indicates the width of the $k$-band used in that data point, and the shaded region indicates the absolute maximum of $k$ we can explore due to our maximum $\ell_{\mathrm{max}}=5000$ choice.
[…]
V Discussion and Conclusions
In Paper I [7], we used the BNT transform to quantify which $\ell$-space truncations in each redshift-bin combination lead to a separation of physical scales in $k$-space. This truncation effectively selects the scale information retained in the cosmological inference; however, Paper I assumed the absence of intrinsic alignments.
Building on the paper I analysis, we examine the behaviour of the BNT transform in two distinct scenarios: one in which the assumed IA model is reasonably accurate but not exact, and another in which it fails to capture the true IA behaviour. We then assess whether the scale separation in $k$-space provided by the BNT method remains valid under these two scenarios.
In the first scenario, the “reasonably accurate” IA case corresponds to using the same IA model for both the sampling and fiducial analyses, while certain fixed parameters of the fiducial model cannot be recovered by the sampling model. We find that the BNT approach still succeeds in separating the physical scales in $k$-space and yields an unbiased estimate of $S_{8}$.
In the second scenario, the fiducial and sampling IA models differ fundamentally—either in redshift or in $k$-space—and can therefore introduce significant biases in the cosmological constraints. The key question is whether we can diagnose inconsistencies in the IA model and thus identify potential biases in the inferred cosmological parameters. Our results show that when the two IA models differ in redshift space, both the noBNT and BNT analyses provide an internal self-consistency check that reveals the discrepancy between the fiducial and sampling models. However, when the IA models differ in $k$-space, only the BNT approach provides the correct diagnostic, the noBNT approach does not provide any means of identifying an incorrect IA model. This is indeed expected as BNT is built to offer a way to separate physical scales in weak lensing.
The main conclusion of this work is that intrinsic alignments do not cause the BNT method to fail. The separation of $k$-scales based on linear combinations of tomographic bins still effectively removes high-$k$ modes where the modelling is unreliable. We also explored the use of $k$-bins as an extension of BNT to isolate physical scales around specific features of interest. Our results show that a reasonably accurate IA model is required but need not be exact, and that when the IA model is fundamentally incorrect, BNT nonetheless provides a consistency test for identifying such cases.
These findings highlight the importance of employing the BNT methodology in Stage-IV lensing analyses to test and mitigate nonlinear and intrinsic-shape systematics, thereby enabling unbiased cosmological constraints with optimal control over scale cuts. This paper serves both as a proof of concept and as an illustration of a general consistency-testing framework. Follow-up studies are anticipated in the following directions:
  • The precision and the characterization of residual systematics of current Stage III cosmic shear surveys is improving, and the first data releases of Stage IV data are also not too far away. The KiDS-Legacy survey [24], with 6 tomographic bins, is a perfect Stage III survey benchmark for testing the BNT transform (Gu et al, in preparation).
  • In paper I, we also proposed a mathematical framework to distribute the constraining power on different parameters on each band of the physical scale. This method still has great potential, like applying consistency checks on the survey setup, or being another method to perform a reconstruction of $P(k)$ with the BNT transform.
Acknowledgements.
All theoretical calculations are based on Pyccl[25] and we present our whole sampling code suit in [29]. We calculate errorbars and plot posterior contours plotting with Getdist[30] package. We are grateful to Hendrik Hildebrandt, Mike Hudson, Bhuvnesh Jain, Benjamin Joachimi, Cora Uhlemann, Angus Wright, Ziang Yan, and the whole German Centre for Cosmological Lensing (GCCL) for their constructive feedback during the construction of the entire structure for the study and the analysis. We thank Sophie Vogt for her help with AxionHMcode. We thank Robert Reischke for providing his covariance matrix calculation code OneCovariance with technical supports.